OpenStudy (anonymous):

What's the product of 4x − 3y)^2??

OpenStudy (diyadiya):

\[(a-b)^2=a^2-2ab+b^2\]

OpenStudy (anonymous):

\[(4x - 3y)^2 = 16x^2 - 24xy + 9y^2\]

OpenStudy (anonymous):

thank you!